OpenStudy (anonymous):

x/4+1=-5 -The / is a fraction. please help!

OpenStudy (kamibug):

Subtract 1 from both sides then multiply both sides by 4.

OpenStudy (anonymous):

so x/4+1-1=-5-1

OpenStudy (kamibug):

You got it, keep going. :)

OpenStudy (anonymous):

1-1=0 -5-1=6

OpenStudy (anonymous):

x/4+0=-6?

OpenStudy (jhannybean):

yes

OpenStudy (kamibug):

Yup. You don't have to write the 0 though. :P x/4=-6. Now multiply both sides by 4. It will cancel on the left, leaving the variable isolated.

OpenStudy (anonymous):

That's how it should look, but you can drop that zero because it's no longer needed. Now to solve for 'x' you need to multiply each side by 4, because you want to isolate the variables.

OpenStudy (anonymous):

Variable(s)

OpenStudy (anonymous):

so x=-24?

OpenStudy (kamibug):

Correct. :)

OpenStudy (anonymous):

Good job.

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

Glad we could help.

OpenStudy (jhannybean):

Don't forget to medal those that helped you the most :)

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

can i medal two people?